# PMGS: Reconstruction of Projectile Motion across Large Spatiotemporal Spans via 3D Gaussian Splatting

### v1

@@ Abstract
Modeling complex rigid motion across large spatiotemporal spans remains an unresolved challenge in dynamic reconstruction. Existing paradigms are mainly confined to short-term, small-scale deformation and offer limited consideration for physical consistency. This study proposes PMGS, focusing on reconstructing Projectile Motion via 3D Gaussian Splatting. The workflow comprises two stages: 1) Target Modeling: achieving object-centralized reconstruction through dynamic scene decomposition and an improved point density control; 2) Motion Recovery: restoring full motion sequences by learning per-frame SE(3) poses. We introduce an acceleration consistency constraint to bridge Newtonian mechanics and pose estimation, and design a dynamic simulated annealing strategy that adaptively schedules learning rates based on motion states. Futhermore, we devise a Kalman fusion scheme to optimize error accumulation from multi-source observations to mitigate disturbances. Experiments show PMGS’s superior performance in reconstructing high-speed nonlinear rigid motion compared to mainstream dynamic methods.
11footnotetext: These authors contributed equally to this work.22footnotetext: * Corresponding authors: Chu He (chuhe@whu.edu.cn), Dingwen Wang (wangdw@whu.edu.cn) and Lei Yu (ly.wd@whu.edu.cn)
1 Introduction
Dynamic reconstruction has become an engine driving modern film animation, game interaction, and virtual reality. The rise of neural rendering has pushed reconstruction fidelity to unprecedented heights, enabling the depiction of highly challenging deformations, such as subtle tremors in biological tissues Endo . Meanwhile, breakthroughs in generative methods Dream4DGS ; Dreamscene4D have achieved controllable synthesis of dynamic scenes with diverse artistic styles, vastly expanding the boundary of imagination in visual expression.
However, returning to a task related to the essential laws of the physical world: reconstruction of complex rigid motions over large spatiotemporal spans—remains an insufficiently explored challenging issue. The difficulties lie in: 1) Modeling paradigm; 2) Physical consistency.
Most existing modeling paradigms are designed for small range non-rigid motion over short time spans DynaNeRF2 ; DyNeRF ; DeformGS ; 4DGS . These approaches rely on deformation fields with complex temporal regularization to achieve spatiotemporal alignment. However, discrete sampling struggles to capture large nonlinear motion (e.g., accelerated rotation), often resulting in noticeable artifacts and trajectory fragmentation when modeling high-speed rigid motion 4DGS . Additionally, current 4D datasets exhibit similar limitations: in most scenarios, moving objects occupy only a small portion of the scene and exhibit minimal movement. Consequently, many methods prioritize optimizing appearance similarity metrics rather than addressing the core challenge of accurate motion recovery.
@@ 1 Introduction
The absence of physical consistency poses another major challenge. Many works focus on optimizing appearance similarity while neglecting the pervasive physical constraints in the real world. However, when modeling complex motion across large spatiotemporal spans, relying solely on photometric supervision fails to constrain the vast 3D solution space effectively. For NeRF-based methods NeRFPlayer ; DyNeRF ; DynaNeRF3 , the inherent characteristics of implicit representations encode motion cues within black-box network weights, making physical interactions difficult to model and unintuitive. The introduction of 3D Gaussian Splatting (3DGS) 3DGS has partially addressed this by combining explicit point-based representations with neural rendering, thereby providing a structured carrier for scene understanding. Consequently, some GS-based approaches Endo ; CF ; PEGAUS incorporate off-the-shelf models to enhance physical realism. We argue that this model-stacking strategy introduces compounded uncertainties by layering new approximations over existing ones. More critically, it bypasses the exploration of fundamental physical principles, ultimately reducing the reconstruction task to uninterpretable compensatory computations.
@@
To tackle these issues, we selected a representative task for focused solutions: Projectile motion. As a fundamental concept in mechanics, it describes an object launched with an initial velocity, following parabolic motion under gravity. The study extends from free-fall to complex situations with self-rotation. This phenomenon is not only ubiquitous in real world (e.g., falling objects from heights, throwing events in sports, military ballistic trajectories), while also encompasses the typical challenges of rigid motion, including variable speed, large span, and compound motion (translation and rotation). Consequently, this archetypal paradigm can be generalized to any dynamic scenario within constant force fields in nature, demonstrating universal applicability.
@@
In this paper, we focus on the modeling of complex rigid motions with large spatiotemporal spans, and propose PMGS, a framework that reconstructs Projectile Motion via 3D Gaussian Splatting. By leveraging temporal continuity in video sequences and explicit Gaussian representation, we estimate per-frame SE(3) affine transformation throughout the motion which fundamentally unattainable with NeRF-based methods. Guided by first principles of physics, we introduce an acceleration constraint for motion recovery. Additionally, to alleviate issues of learning oscillations and trajectory fractures caused by fixed training paradigms, we develop a Dynamic Simulated Annealing (DSA) strategy that adaptively schedules the training process based on real-time velocity and displacement variations. Departing from conventional methods that treat physical constraints merely as regularization components, we propose a Kalman filter-based optimal estimation scheme to fuse cross-modal observations thereby minimizing potential disturbances.
@@
  • We design a Kalman fusion scheme is to dynamically optimize multi-source observations with adaptive weighting, minimizing error accumulation caused by real-world disturbances or training oscillations.
@@
Figure 2: Overview of PMGS. We first segment the target then decompose the motion through centralization to transform the dynamic scene into static. For modeling, we learn a set of Gaussian kernels and align them at the original scale with a set of learnable affine transformation. In the motion recovery, we estimate the target’s SE(3) transformation frame by frame, and comprehensively improve tracking accuracy by integrating physics-enhanced strategies.
2 Related Work
2.1 3D Representation.
Im1 ; Im3 pioneered the exploration of the implicit radiance fields.The milestone work, NeRF NeRF , maps spatial coordinates and viewing directions to color and density through the MLP, enabling differentiable and compact units of complex scenes. However, the implicit representations are inherently constrained in geometric nature, which limits their exploration for editing and interaction. Instead, 3DGS 3DGS introduces an shift from implicit to explicit paradigm, the point-based representation greatly expands its potential of scene understanding.
2.2 Dynamic Neural Rendering.
@@ 2 Related Work
Current dynamic neural rendering include: (1) Point deformation fields DynaNeRF2 map sampled points at each timestep to a static canonical space. (2) Time-aware Volume Rendering DyNeRF ; NeRFPlayer employs factorized voxels to independently compute features for each point. (3) The Gaussian deformation field DeformGS ; MotionGS ; S4d transforms 3D Gaussians to their target position at specific timesteps. The critical issues lie in the deformation field struggles to model large inter-frame displacements that commonly occur in real-world scenarios, which can be further validated by experiments on Panoptic Studio dataset Sportsdatasest conducted by 4DGS 4DGS . Moreover, due to the lack of physical consistency constraints, the network tends to focus on fitting appearance similarity while neglecting structural accuracy at the geometric level. As a typical example, when there are large areas of nearly uniform color, Gaussian bodies will float chaotically within regions of similar color Dynamic3DGS .
2.3 Enhancing GS with Physical Assistance.
@@
Introducing physical assistance to enhance 3DGS has been widely applied. CF ; Endo ; FSGS utilize off-the-shelf monocular depth estimation models DPT for scene initialization or sparse reconstruction. Spring devised a Spring-Mass model to simulate the elastic objects’s falling and collision. PEGAUS is designed for 3D dataset generation which integrates the PyBullet engine to emulate the placement of objects and their dynamic processes. Dreamscene4D ; SAGS enhances pose estimation by ensuring scale consistency and local rigidity. Essentially, current approaches use rough approximations from external physics models to fix ill-posed inverse problems in vision systems. Nevertheless, pre-trained engines lose confidence sharply in unusual scenarios, and their black-box makes errors untraceable—causing mistakes to snowball. Differently, we propose to explore strategies grounded directly in first principles of physics particlegs , rather than relying on external models for unreliable assistance.
3 Method
@@
PMGS takes a monocular video as input to reconstruct the target and achieve full-sequence projectile motion recovery. For modeling appearance and geometry, we equivalently converts dynamic scenes into static through motion decomposition, and combined with the improved point density control strategy to enhance the geometric accuracy. During motion recovery, we leverage video temporal continuity and the explicit Gaussian representation to estimate per-frame SE(3) transformations. To ensure physically consistent pose learning, we introduce an acceleration consistency constraint that establishes direct connections between pose estimation and Newtonian dynamics priors. Furthermore, we devise a DSA strategy, which adaptively schedules the optimization process according to real-time motion states. Finally, we design a Kalman fusion strategy to optimize the error accumulation across multi-modal observation sources.
3.1 Target modeling
@@
Centralization. We firstly use the pre-trained SAM SAM to separate the background and the dynamic target. Then, by establishing an object-centralized normalized coordinate Dreamscene4D , we decompose the complex projectile motion: the object’s autorotation is disentangled into pseudo multi-view observations, while translational displacements are eliminated.
Reconstruction. Following the 3DGS 3DGS , we use a set of Gaussian kernels to explicitly reconstruct the appearance and geometry of the object.
@@
where $\Sigma_{i}$ and $\mu_{i}$ represent the covariance matrix and position of a Gaussian $g_{i}$ respectively. For axial constraint, Gaussians with length of principal axis $\max\big{(}\mathrm{eig}(\Sigma_{i})\big{)}>\tau_{L}$ are discarded, where $\tau_{L}$ is a predefined threshold controlling shape anisotropy. Besides, an outlier is defined as a Gaussian whose minimum distance to other is much larger than the average distance $D_{avg}=\frac{1}{|\mathcal{G}|^{2}}\sum_{m,n}\|\mu_{m}-\mu_{n}\|_{2}$, where $\tau_{D}$ is a distance-based filtering factor.
During the optimization, the loss $\mathcal{L}_{GS}$ is used to quantify the difference between the real image $\hat{I}$ and render image $I$:
@@
Up to this point, we consequently obtain a Gaussian field $G_{0}$ at the centralized scale.
Registration. Finally, we register the static Gaussian $G_{0}$ at the centralized scale to the original dynamic scene, thus facilitating subsequent motion recovery. This can be achieved via a set of learnable affine transformations $T_{reg}=[\mathbf{r,t},s]$, i.e., quaternion rotation $\mathbf{r}\in\mathbb{R}^{4}$, translation vector $\mathbf{t}\in\mathbb{R}^{3}$, and constant scalar scale $s$. Then the aligned Gaussian field $G_{1}$ can be represented as $T_{reg}\odot G_{0}$. In addition to $\mathcal{L}_{GS}$, we further apply VGG loss $\mathcal{L}_{VGG}$ and grid-based total variation loss $\mathcal{L}_{TV}$ Spring ; 4DGS to achieve a better awareness for accurate registration. The total loss can be formulated as:
@@
Finally, we obtain this set of Gaussian $G_{1}$, which can well-represent the appearance and geometry of the object.
Figure 3: DSA strategy. Right: An object accelerates in a constant gravity field, with different velocity $v$ and displacements $s$ corresponding to each timestamp. Left: Red curve represents the initial learning rate. Blue points shows the final learning rate after exponential decay.
3.2 Physics-Enhanced Motion recovery
During the dynamic phase, we estimate 6DoF pose changes of the object in total $(n+1)$ frames, which corresponds to a sequence of temporal SE-3 transformations $T_{n}=[\mathbf{r}_{n},\mathbf{t}_{n}]$. We introduce an acceleration consistency constraint to establish a direct connection between pose learning and physical priors. Additionally, we incorporate optical flow smoothing and a DSA training strategy to improve training stability.
Acceleration consistency constraint. The object moves in a constant force field, therefore its acceleration remains constant. Based on this, we first calculate the object’s center of mass $\sigma$. Since the Gaussian have been processed isotropically, the spatial coordinates of $\sigma$ can be computed as:
@@ 3.2 Physics-Enhanced Motion recovery
According to the Newtonian dynamics, there should be a constant acceleration along the gravity direction, while the components orthogonal to it should remain zero:
$\mathcal{L}_{Acc}=\lVert{\mathbf{a}^{t}_{\parallel}+(a_{\perp}^{(t+\Delta{t})}-a_{\perp}^{t})}\rVert_{2}^{2}$ (8)
@@
Dynamic Simulated Annealing. As shown in Figure 3, a fixed $lr_{init}$ struggles to adapt to varying object’s velocities: excessively large $lr_{init}$ induce training oscillations during low-speed phases, whereas small $lr_{init}$ hinder convergence when tracking high-speed targets. To address this velocity-displacement coupling effect, we dynamically schedule the $lr_{init}$ according to real-time displacement.
@@
As the initial velocity $v_{0}$, acceleration $a$ and the time interval $\Delta t$ are constant value, thus Equation 9 establishes the linear correlation between object’s displacement $\Delta s$ and timestamp $t$, i.e., $\Delta s\propto t$. This temporal-displacement relationship therefore requires proportional scaling of the $lr_{init}$ with displacement. Therefore, we implement scheduling where the learned displacement from the preceding timestep governs the subsequent $lr_{init}$, as visualized in Figure 3.
As the iteration progresses, the Gaussians gradually approach the target spatial position. Correspondingly, the learning rate needs to be reduced so that the step gradually shrinks for finetuning. We utilized the commonly used exponential decay to achieve this. In addition, for timestamps with large displacements, we appropriately extend the number of iterations to better find the optimal solution.
@@
$T_{n}$ $=\mathrm{arg}\underset{T_{n}}{\mathrm{min}}(\lambda_{1}\mathcal{L}_{GS}(\mathcal{R}(G_{n}),\hat{I}_{n})+\lambda_{2}\mathcal{L}_{Acc})$ (10)
3.3 Cross-modal Kalman fusion
Furthermore, we introduce optical flow smoothing computed based on Optical for emphasizing attention to motion variations in long-term tracking. Though some works MotionAware also incorporate optical flow or priors as auxiliaries, but they fail to account for the inaccurate optical flow estimation caused by degraded textures under motion blur, which in turn introduces negative effects. Differently, we design a data fusion strategy based on Kalman Filter Kalman to dynamically balance the weights of observation sources, suppress cumulative error, and update the optimal estimate in real-time.
@@ 3.3 Cross-modal Kalman fusion
$\textbf{X}_{t|t-1}=\begin{bmatrix}\Delta s_{t|t-1}\\ v_{t|t-1}\end{bmatrix}=\textbf{F}\textbf{X}_{t-1|t-1}+\textbf{B}a_{k}+\textbf{w}_{k},$ (11)
@@
where $\textbf{X}_{t|t-1}$ is the prediction based on the optimal estimate at previous timestamp, F is the state transition matrix, B is the control input matrix, and $a_{t}$ is the control input calculated from Eq.6, which approximates to $a_{t-1}$ as the system model defined. $\textbf{Q}=\mathbb{E}[\textbf{w}_{k}\textbf{w}_{k}^{T}]$ is the covariance matrix of the process noise $\textbf{w}_{k}$ (includes noise $\sigma_{\Delta s}^{2}$ and $\sigma_{v}^{2}$) that describes the uncertainty of the model. The prediction covariance then can be calculated as:
@@
Step4-Update: Based on above content, we update the Kalman gain $\textbf{K}_{t}$ and the posterior covariance $\textbf{P}_{t|t}$:

### v3

@@ Abstract
Modeling complex rigid motion across large spatiotemporal spans remains an unresolved challenge in dynamic reconstruction. Existing paradigms are mainly confined to short-term, small-scale deformation and offer limited consideration for physical consistency. This study proposes PMGS, focusing on reconstructing Projectile Motion via 3D Gaussian Splatting. The workflow comprises two stages: 1) Target Modeling: achieving object-centralized reconstruction through dynamic scene decomposition and an improved point density control; 2) Motion Recovery: restoring full motion sequences by learning per-frame SE-3 poses. We introduce an acceleration consistency constraint to bridge Newtonian mechanics and pose estimation, and design a dynamic simulated annealing strategy that adaptively schedules learning rates based on motion states. Furthermore, we devise a Kalman fusion scheme to optimize error accumulation from multi-source observations to mitigate disturbances. Experiments show PMGS’s superior performance in reconstructing high-speed nonlinear rigid motion compared to mainstream dynamic methods.
Codehttps://github.com/X-Probiotics/PMGS
Introduction
Dynamic reconstruction has become an engine driving modern film animation, game interaction, and virtual reality. The rise of neural rendering has pushed reconstruction fidelity to unprecedented heights, enabling the depiction of highly challenging deformations, such as subtle tremors in biological tissues (Endo). Meanwhile, breakthroughs in generative methods (Dream4DGS; Dreamscene4D) have achieved controllable synthesis of dynamic scenes with diverse artistic styles, vastly expanding the boundary of imagination in visual expression.
However, when returning to a task governed by the essential laws of the physical world—the reconstruction of complex rigid motions over large spatiotemporal spans—there lies a challenging issue that remains insufficiently explored.
Most existing modeling paradigms are designed for small range non-rigid motion over short time spans (DynaNeRF2; DyNeRF; DeformGS; 4DGS). These approaches rely on deformation fields with complex temporal regularization to achieve spatiotemporal alignment. However, discrete sampling struggles to capture large nonlinear motion (e.g., accelerated rotation), often resulting in noticeable artifacts and trajectory fragmentation when modeling high-speed rigid motion (4DGS). Additionally, current 4D datasets exhibit similar limitations: in most scenarios, moving objects occupy only a small portion of the scene and exhibit minimal movement. Consequently, many methods prioritize optimizing appearance similarity metrics rather than addressing the core challenge of accurate motion recovery.
@@ Introduction
The absence of physical consistency poses another major challenge. Many works focus on optimizing appearance similarity while neglecting the pervasive physical constraints in the real world. However, when modeling complex motion across large spatiotemporal spans, relying solely on photometric supervision fails to constrain the vast 3D solution space effectively. For NeRF-based methods (NeRFPlayer; DyNeRF; DynaNeRF3), the inherent characteristics of implicit representations encode motion cues within black-box network weights, making physical interactions difficult to model and unintuitive. The introduction of 3D Gaussian Splatting (3DGS) (3DGS) has partially addressed this by combining explicit point-based representations with neural rendering, thereby providing a structured carrier for scene understanding. Consequently, some GS-based approaches (Endo; CF; PEGAUS) incorporate off-the-shelf models to enhance physical realism. We argue that this model-stacking strategy introduces compounded uncertainties by layering new approximations over existing ones. More critically, it bypasses the exploration of fundamental physical principles, ultimately reducing the reconstruction task to uninterpretable compensatory computations.
@@
To tackle these issues, we selected a representative task for focused solutions: Projectile motion. As a fundamental concept in mechanics, it describes an object launched with an initial velocity, following parabolic motion under gravity. The study extends from free-fall to complex situations with self-rotation. This phenomenon is not only ubiquitous in the real world (e.g., falling objects from heights, throwing events in sports, military ballistic trajectories), but also encompasses the typical challenges of rigid motion, including variable speed, large span, and compound motion (translation and rotation). Consequently, this archetypal paradigm can be generalized to any dynamic scenario within constant force fields in nature, demonstrating universal applicability.
@@
In this paper, we focus on the modeling of complex rigid motions with large spatiotemporal spans, and propose PMGS, a framework that reconstructs Projectile Motion via 3D Gaussian Splatting. By leveraging temporal continuity in video sequences and explicit Gaussian representation, we estimate per-frame SE(3) affine transformation throughout the motion which is fundamentally unattainable with NeRF-based methods. Guided by first principles of physics, we introduce an acceleration constraint for motion recovery. Additionally, to alleviate issues of learning oscillations and trajectory fractures caused by fixed training paradigms, we develop a Dynamic Simulated Annealing (DSA) strategy that adaptively schedules the training process based on real-time velocity and displacement variations. Departing from conventional methods that treat physical constraints merely as regularization components, we propose a Kalman filter-based optimal estimation scheme to fuse cross-modal observations, thereby minimizing potential disturbances.
@@
  • We design a Kalman fusion scheme to dynamically optimize multi-source observations with adaptive weighting, minimizing error accumulation caused by real-world disturbances or training oscillations.
@@
Figure 2: Overview of PMGS. We first segment the target, then decompose the motion through centralization to transform the dynamic scene into a static one. For modeling, we learn a set of Gaussian kernels and align them at the original scale with a set of learnable affine transformations. In motion recovery, we estimate the target’s SE(3) transformation frame by frame, and comprehensively improve tracking accuracy by integrating physics-enhanced strategies.
Related Work
Dynamic Neural Rendering.
@@
Current dynamic neural rendering includes: (1) Point deformation fields (DynaNeRF2) map sampled points at each timestep to a static canonical space. (2) Time-aware Volume Rendering (DyNeRF; NeRFPlayer) employs factorized voxels to independently compute features for each point. (3) The Gaussian deformation field (DeformGS; MotionGS; S4d) transforms 3D Gaussians to their target position at specific timesteps. The critical issues lie in that the deformation field struggles to model large inter-frame displacements that commonly occur in real-world scenarios, which can be further validated by experiments on Panoptic Studio dataset (Sportsdatasest) conducted by 4DGS (4DGS). Moreover, due to the lack of physical consistency constraints, the network tends to focus on fitting appearance similarity while neglecting structural accuracy at the geometric level. As a typical example, when there are large areas of nearly uniform color, Gaussian bodies will float chaotically within regions of similar color (Dynamic3DGS).
Enhancing GS with Physical Assistance.
@@
Introducing physical assistance to enhance 3DGS has been widely applied. (CF; Endo; FSGS) utilize off-the-shelf monocular depth estimation models (DPT) for scene initialization or sparse reconstruction. (Spring) devised a Spring-Mass model to simulate the elastic objects’ falling and collision. (PEGAUS) is designed for 3D dataset generation which integrates the PyBullet engine to emulate the placement of objects and their dynamic processes. (Dreamscene4D; SAGS) enhances pose estimation by ensuring scale consistency and local rigidity. Essentially, current approaches use rough approximations from external physics models to fix ill-posed inverse problems in vision systems. Nevertheless, pre-trained engines lose confidence sharply in unusual scenarios, and their black-box makes errors untraceable.
Method
@@
PMGS takes a monocular video as input to reconstruct the target and achieve full-sequence projectile motion recovery. For modeling appearance and geometry, we equivalently convert dynamic scenes into static through motion decomposition, and combine with the improved point density control strategy to enhance the geometric accuracy. During motion recovery, we leverage video temporal continuity and the explicit Gaussian representation to estimate per-frame SE(3) transformations. To ensure physically consistent pose learning, we introduce an acceleration consistency constraint that establishes direct connections between pose estimation and Newtonian dynamics priors. Finally, we design a Kalman fusion strategy to optimize the error accumulation from multi-modal observation sources.
Target Modeling
@@
Centralization. We firstly use the pre-trained SAM (SAM) to separate the background and the dynamic target. Then, by establishing object-centralized normalized coordinates (Dreamscene4D), we decompose the complex projectile motion: the object’s autorotation is disentangled into pseudo multi-view observations, while translational displacements are eliminated.
Reconstruction. Following the 3DGS (3DGS), we use a set of Gaussian kernels to explicitly reconstruct the appearance and geometry of the object.
@@
where $\Sigma_{i}$ and $\mu_{i}$ represent the covariance matrix and position of a Gaussian $g_{i}$ respectively. For axial constraint, Gaussians with length of principal axis $\max\big(\mathrm{eig}(\Sigma_{i})\big)>\tau_{L}$ are discarded, where $\tau_{L}$ is a predefined threshold controlling shape anisotropy. Besides, an outlier is defined as a Gaussian whose minimum distance to others is much larger than the average distance $D_{avg}=\frac{1}{|\mathcal{G}|^{2}}\sum_{m,n}\|\mu_{m}-\mu_{n}\|_{2}$, where $\tau_{D}$ is a distance-based filtering factor.
During the optimization, the loss $\mathcal{L}_{GS}$ is used to quantify the difference between the real image $\hat{I}$ and rendered image $I$:
@@
Up to this point, we obtain a Gaussian field $G_{0}$ at the centralized scale.
Registration. Finally, we register the static Gaussian $G_{0}$ at the centralized scale to the original dynamic scene, thus facilitating subsequent motion recovery. This can be achieved via a set of learnable affine transformations $T_{reg}=[\mathbf{r,t},s]$. Then the aligned Gaussian field $G_{1}$ can be represented as $T_{reg}\odot G_{0}$. In addition to $\mathcal{L}_{GS}$, we further apply VGG loss $\mathcal{L}_{VGG}$ and grid-based total variation loss $\mathcal{L}_{TV}$ (Spring; 4DGS) to achieve a better awareness of accurate registration. The total loss can be formulated as:
@@
Finally, we obtain this set of Gaussians $G_{1}$, which can well represent the appearance and geometry of the object.
Figure 3: DSA strategy. Right: An object accelerates in a constant gravity field, with different velocities $v$ and displacements $s$ corresponding to each timestamp. Left: Red curve represents the initial learning rate. Blue points show the final learning rate after exponential decay.
Physics-Enhanced Motion Recovery
During the dynamic phase, we estimate 6DoF pose changes of the object in total $(n+1)$ frames, which correspond to a sequence of temporal SE-3 transformations $T_{n}=[\mathbf{r}_{n},\mathbf{t}_{n}]$. We introduce an acceleration consistency constraint to establish a direct connection between pose learning and physical priors. Additionally, we incorporate optical flow smoothing and a DSA training strategy to improve training stability.
Acceleration consistency constraint. The object moves in a constant force field, therefore its acceleration remains constant. Based on this, we first calculate the object’s center of mass $\sigma$. Since the Gaussians have been processed isotropically, the spatial coordinates of $\sigma$ can be computed as:
@@
According to Newtonian dynamics, there should be a constant acceleration along the gravity direction, while the components orthogonal to it should remain zero:
$\mathcal{L}_{Acc}=\lVert{\mathbf{a}^{t}_{\parallel}+(\mathbf{a}_{\perp}^{(t+\Delta{t})}-\mathbf{a}_{\perp}^{t})}\rVert_{2}^{2}$ (8)
@@
Optical flow smoothing. We introduce optical flow smoothing computed based on (Optical) for emphasizing attention to motion variations in long-term tracking. In addition to evaluating the similarity between the real optical flow $\hat{F}$ and rendered $F$, we compute the gradients of the horizontal and vertical components of the optical flow field, applying a smoothness penalty with different weights for corresponding regions:
$\mathcal{L}_{Smooth}$ $=\lambda_{1}[\frac{1}{N}\sum_{i=1}^{N}(\Delta F_{i,x}\cdot\exp(-\frac{\Delta I_{i,x}}{10})+$ (9)
$\Delta F_{i,y}\cdot\exp(-\frac{\Delta I_{i,y}}{10}))]+\lambda_{2}\lVert{\hat{F}-F}\rVert_{1}$
where $\Delta F$ and $\Delta I$ denote the gradients of the optical flow field and the rendered image respectively.
Dynamic simulated annealing. As shown in Figure 3, a fixed $lr_{init}$ struggles to adapt to varying object’s velocities: excessively large $lr_{init}$ induces training oscillations during low-speed phases, whereas small $lr_{init}$ hinders convergence when tracking high-speed targets. To address this velocity-displacement coupling effect, we dynamically schedule the $lr_{init}$ according to real-time displacement.
@@
As the initial velocity $v_{0}$, acceleration $a$ and the time interval $\Delta t$ are constant values, Eq.10 establishes the linear correlation between object’s displacement $\Delta s$ and timestamp $t$, i.e., $\Delta s\propto t$. This temporal-displacement relationship therefore requires proportional scaling of the $lr_{init}$ with displacement. Therefore, we implement scheduling where the learned displacement from the preceding timestep governs the subsequent $lr_{init}$, as visualized in Figure 3.
As the iteration progresses, the Gaussians gradually approach the target spatial position. Correspondingly, the learning rate needs to be reduced so that the step gradually shrinks for finetuning. We utilize the commonly used exponential decay to achieve this. In addition, for timestamps with large displacements, we appropriately extend the number of iterations to better find the optimal solution.
@@
$T_{n}$ $=\mathrm{arg}\underset{T_{n}}{\mathrm{min}}(\lambda_{1}\mathcal{L}_{GS}(\mathcal{R}(T_{n}\odot G_{n}),\hat{I}_{n+1})+$ (11)
$\lambda_{2}\mathcal{L}_{Acc}+\lambda_{3}\mathcal{L}_{Smooth}({F}(T_{n}\odot G_{n}),\hat{F}_{n+1})$
Cross-Modal Kalman Fusion
During the aforementioned stage, there exist different sources of cross-modal observations. Some works (MotionAware) also incorporate optical flow or other priors as auxiliary inputs. However, they often fail to account for inaccurate optical flow estimations caused by motion-blurred and texture-degraded regions. This oversight, in turn, introduces negative effects during optimization. Differently, we design a data fusion strategy based on Kalman filter (Kalman) to dynamically balance the weights of observation sources, suppress cumulative error, and update the optimal estimate in real-time.
@@
$\textbf{X}_{t|t-1}=\begin{bmatrix}\Delta s_{t|t-1}\\ v_{t|t-1}\end{bmatrix}=\textbf{F}\textbf{X}_{t-1|t-1}+\textbf{B}a_{t}+\textbf{w}_{t},$ (12)
@@
where $\textbf{X}_{t|t-1}$ is the prediction based on the optimal estimate at the previous timestamp, F is the state transition matrix, B is the control input matrix, and $a_{t}$ is the control input calculated from Eq.6, which approximates to $a_{t-1}$ as the system model defined. The prediction covariance can then be calculated as:
@@
Datasets Source Type Phy. Mono.
LaSOT (lasot) Real Rigid $\times$ $\times$
D-NeRF (DynaNeRF2) Syn. Def. $\times$ $\checkmark$
HyperNeRF (DynaNeRF3) Real Def. $\times$ $\checkmark$
Neural3D (DyNeRF) Syn. Def. $\times$ $\times$
PEGASUS (PEGAUS) Syn. Rigid PyBullet $\checkmark$
SpringGS (Spring) 6S + 3R Def. Hooke’s Law $\times$
PMGS 6S + 3R Rigid Newton’s Law $\checkmark$
Grenade Missile Cola Doll Chocolate Rugby
@@
2DGS 24.82 0.895 0.153 17.42 0.714 0.325 15.97 0.847 0.195 17.99 0.792 0.218 15.19 0.758 0.275 21.56 0.830 0.185
In-Splat 25.44 0.900 0.139 17.58 0.724 0.300 17.03 0.852 0.175 18.10 0.797 0.221 16.10 0.761 0.255 24.21 0.877 0.165
CFGS* 21.82 0.873 0.231 13.70 0.675 0.259 14.47 0.845 0.284 18.00 0.788 0.307 11.93 0.739 0.317 21.37 0.822 0.196
4DGS 20.52 0.907 0.112 22.60 0.829 0.082 20.25 0.859 0.154 23.92 0.947 0.063 17.91 0.934 0.182 Failed Failed Failed
Mo-GS 22.64 0.876 0.166 19.14 0.857 0.156 15.39 0.823 0.217 22.70 0.854 0.119 16.99 0.902 0.115 17.37 0.855 0.151
Dy-GS 36.32 0.971 0.064 26.37 0.906 0.032 28.44 0.939 0.026 31.81 0.942 0.050 25.36 0.897 0.018 Failed Failed Failed
PMGS 33.58 0.967 0.011 28.87 0.915 0.025 26.50 0.940 0.014 34.38 0.959 0.014 25.01 0.924 0.020 33.72 0.954 0.012
Table 2: Comparison of video reconstruction (Synthetic data).
Box Shark Bear
[…]
2DGS 19.72 0.973 0.172 11.57 0.794 0.101 11.78 0.823 0.201
In-Splat 20.48 0.977 0.152 12.88 0.810 0.097 11.79 0.823 0.090
CFGS* 17.19 0.946 0.162 11.32 0.742 0.271 11.54 0.772 0.212
4DGS 20.19 0.874 0.131 15.32 0.709 0.302 Failed Failed Failed
Mo-GS 15.92 0.512 0.464 10.38 0.654 0.353 13.99 0.887 0.231
Dy-GS 18.97 0.704 0.152 24.92 0.952 0.085 Failed Failed Failed
PMGS 30.45 0.974 0.019 25.07 0.952 0.094 26.64 0.959 0.059
Table 3: Comparison of video reconstruction (Real data).
Grenade Missile Cola Doll Chocolate Rugby
[…]
CFGS* 0.249 0.550 0.602 0.203 0.129 0.218 0.154 0.533 0.563 0.310 0.358 0.459 0.318 0.574 0.610 0.094 0.562 0.604
4DGS 0.359 0.661 0.703 0.608 0.194 0.248 0.354 0.630 0.661 0.401 0.563 0.621 0.287 0.668 0.687 Failed Failed Failed
Mo-GS 0.288 0.474 0.505 0.541 0.655 0.704 0.423 0.511 0.551 0.501 0.505 0.559 0.549 0.501 0.522 0.249 0.643 0.668
Dy-GS 0.548 0.227 0.269 0.910 0.118 0.156 0.898 0.182 0.226 0.663 0.311 0.363 0.995 0.070 0.090 Failed Failed Failed
PMGS 0.998 0.168 0.254 0.995 0.091 0.130 0.997 0.069 0.080 0.993 0.092 0.152 0.987 0.168 0.214 0.952 0.021 0.148
Table 4: Comparison of motion recovery (Synthetic data).
Box Shark Bear
[…]
CFGS* 0.079 0.392 0.425 0.252 0.425 0.458 0.199 0.319 0.364
4DGS 0.144 0.553 0.592 0.141 0.505 0.556 Failed Failed Failed
Mo-GS 0.299 0.267 0.313 0.298 0.557 0.629 0.179 0.526 0.573
Dy-GS 0.262 0.439 0.470 0.679 0.094 0.155 Failed Failed Failed
PMGS 0.940 0.055 0.086 0.876 0.021 0.093 0.969 0.043 0.047
Table 5: Comparison of motion recovery (Real data).
Video reconstruction Motion recovery
[…]
CFGS* 16.82 0.799 0.249 0.206 0.426 0.478
Model 1 24.61 0.910 0.089 0.635 0.334 0.791
Model 2 25.63 0.929 0.045 0.886 0.137 0.188
Model 3 28.95 0.940 0.035 0.933 0.098 0.147
Full 29.35 0.949 0.029 0.967 0.080 0.133
Table 6: Ablation study of different models.
Step3-Correction: In the update step, we use observations (optical flow and learned displacement) to correct the predicted state. These two measurements provide direct observations of displacement, but each carries distinct noise: (1) $z_{flow}^{t}$ is computed from optical flow (Eq.9), representing apparent motion-based displacement observation, with noise denoted as $\sigma_{flow}^{2}$. (2) $z_{learn}^{t}$ derives from the optimization process (Eq. 11), based on Gaussian rendering and loss functions, with noise denoted as $\sigma_{learn}^{2}$. Thus, we define the observation equation as follows:
@@
Step4-Update: Based on the above content, we update the Kalman gain $\textbf{K}_{t}$: